\begin{thebibliography}{}

\bibitem[\protect\citename{Ammar \bgroup et al.\egroup
  }2014]{ammar2014conditional}
Waleed Ammar, Chris Dyer, and Noah~A Smith.
\newblock 2014.
\newblock Conditional random field autoencoders for unsupervised structured
  prediction.
\newblock In {\em Advances in Neural Information Processing Systems}, pages
  3311--3319.

\bibitem[\protect\citename{Bastien \bgroup et al.\egroup }2012]{Bastien:2012}
Frederic Bastien, Pascal Lamblin, Razvan Pascanu, James Bergstra, Ian
  Goodfellow, Arnaud Bergeron, Nicolas Bouchard, David Warde-Farley, and Yoshua
  Bengio.
\newblock 2012.
\newblock {Theano: new features and speed improvements}.
\newblock In {\em NIPS 2012 deep learning workshop}.

\bibitem[\protect\citename{Bengio \bgroup et al.\egroup }1994]{Yoshua:1994}
Yoshua Bengio, Patrice Simard, and Paolo Frasconi.
\newblock 1994.
\newblock {Learning long-term dependencies with gradient descent is difficult}.
\newblock In {\em IEEE Transactions on Neural Networks}, pages 157--166.

\bibitem[\protect\citename{Bengio \bgroup et al.\egroup }2003]{Bengio:2003}
Yoshua Bengio, Rejean Ducharme, Pascal Vincent, and Christian Jauvin.
\newblock 2003.
\newblock {A neural probabilistic language model}.
\newblock In {\em Journal of Machine Learning Reseach}, volume~3.

\bibitem[\protect\citename{Bengio}2009]{bengio:2009}
Yoshua Bengio.
\newblock 2009.
\newblock {Learning Deep Architectures for AI}.
\newblock {\em Foundations and trends{\textregistered} in Machine Learning},
  2(1):1--127.

\bibitem[\protect\citename{Bergstra \bgroup et al.\egroup }2010]{Bergstra:2010}
James Bergstra, Olivier Breuleux, Frederic Bastien, Pascal Lamblin, Razvan
  Pascanu, Guillaume Desjardins, Joseph Turian, David Warde-Farley, and Yoshua
  Bengio.
\newblock 2010.
\newblock {Theano: a CPU and GPU math expression compiler}.
\newblock In {\em Proceedings of the Python for scientific computing conference
  (SciPy)}, volume~4.

\bibitem[\protect\citename{Cho \bgroup et al.\egroup }2014]{Cho:2014}
Kyunghyun Cho, Bart~Van Merriënboer, Caglar Gulcehre, Dzmitry Bahdanau, Fethi
  Bougares, Holger Schwenk, and Yoshua Bengio.
\newblock 2014.
\newblock {Learning phrase representations using RNN encoder-decoder for
  statistical machine translation}.
\newblock In {\em Proceedings of the Conference on Empirical Methods in Natural
  Language Processing (EMNLP)}, Doha, Qatar.

\bibitem[\protect\citename{Chung \bgroup et al.\egroup }2014]{Chung:2014}
Junyoung Chung, Caglar Gulcehre, KyungHyun Cho, and Yoshua Bengio.
\newblock 2014.
\newblock {Empirical evaluation of gated recurrent neural networks on sequence
  modeling}.
\newblock In {\em arXiv:1412.3555 [cs.NE]}.

\bibitem[\protect\citename{Das and Petrov}2011]{das:2011}
Dipanjan Das and Slav Petrov.
\newblock 2011.
\newblock {Unsupervised part-of-speech tagging with bilingual graph-based
  projections}.
\newblock In {\em Proceedings of the 49th Annual Meeting of the Association for
  Computational Linguistics: Human Language Technologies-Volume 1}, pages
  600--609. Association for Computational Linguistics.

\bibitem[\protect\citename{Duong \bgroup et al.\egroup }2013]{duong:2013}
Long Duong, Paul Cook, Steven Bird, and Pavel Pecina.
\newblock 2013.
\newblock Simpler unsupervised pos tagging with bilingual projections.
\newblock In {\em ACL (2)}, pages 634--639.

\bibitem[\protect\citename{Elman}1990]{elman:1990}
Jeffrey~L Elman.
\newblock 1990.
\newblock {Finding Structure in Time}.
\newblock {\em Cognitive science}, 14(2):179--211.

\bibitem[\protect\citename{Erhan \bgroup et al.\egroup }2010]{erhan:2010}
Dumitru Erhan, Yoshua Bengio, Aaron Courville, Pierre-Antoine Manzagol, Pascal
  Vincent, and Samy Bengio.
\newblock 2010.
\newblock {Why Does Unsupervised Pre-training Help Deep Learning?}
\newblock {\em Journal of Machine Learning Research}, 11(Feb):625--660.

\bibitem[\protect\citename{Gouws and S{\o}gaard}2015]{gouws:2015}
Stephan Gouws and Anders S{\o}gaard.
\newblock 2015.
\newblock Simple task-specific bilingual word embeddings.
\newblock In {\em Proceedings of NAACL-HLT}, pages 1386--1390.

\bibitem[\protect\citename{Graves}2013]{Graves:2013}
Alex Graves.
\newblock 2013.
\newblock {Generating sequences with recurrent neural networks}.
\newblock {\em arXiv:1308.0850}.

\bibitem[\protect\citename{Hal{\'a}csy \bgroup et al.\egroup
  }2007]{hunpos:2007}
P{\'e}ter Hal{\'a}csy, Andr{\'a}s Kornai, and Csaba Oravecz.
\newblock 2007.
\newblock Hunpos: An open source trigram tagger.
\newblock In {\em Proceedings of the 45th annual meeting of the ACL on
  interactive poster and demonstration sessions}, pages 209--212. Association
  for Computational Linguistics.

\bibitem[\protect\citename{Hochreiter and Schmidhuber}1997]{Hochreiter:1997}
Sepp Hochreiter and Jurgen Schmidhuber.
\newblock 1997.
\newblock {Long short-term memory}.
\newblock In {\em Neural computation}, pages 1735--1780.

\bibitem[\protect\citename{Jamatia and Das}2016]{jamatia:2016}
Anupam Jamatia and Amitava Das.
\newblock 2016.
\newblock {Task Report: Tool Contest on POS Tagging for Code-Mixed Indian
  Social Media (Facebook, Twitter, and Whatsapp) Text@ICON 2016}.
\newblock In {\em Proceedings of ICON 2016}.

\bibitem[\protect\citename{Jamatia \bgroup et al.\egroup }2015]{jamatia:2015}
Anupam Jamatia, Bj{\"o}rn Gamb{\"a}ck, and Amitava Das.
\newblock 2015.
\newblock {Part-of-speech tagging for code-mixed English-Hindi Twitter and
  Facebook chat messages}.
\newblock {\em RECENT ADVANCES IN}, page 239.

\bibitem[\protect\citename{Jordan}1986]{jordan:1986}
Michael~I Jordan.
\newblock 1986.
\newblock {Attractor Dynamics and Parallellism in a Connectionist Sequential
  Machine}.
\newblock In {\em Proceedings of 1986 Cognitive Science Conference}, pages
  531--546.

\bibitem[\protect\citename{Jozefowicz \bgroup et al.\egroup
  }2015]{Jozefowicz:2015}
Rafal Jozefowicz, Wojciech Zaremba, and Ilya Sutskever.
\newblock 2015.
\newblock {An empirical exploration of recurrent network architectures}.
\newblock In {\em Proceedings of the 32nd International Conference on Machine
  Learning}, pages 2342--2350.

\bibitem[\protect\citename{Kreutzer \bgroup et al.\egroup }2015]{kreutzer:2015}
Julia Kreutzer, Shigehiko Schamoni, and Stefan Riezler.
\newblock 2015.
\newblock {QUality Estimation from ScraTCH (QUETCH): Deep Learning for
  Word-level Translation Quality Estimation}.
\newblock In {\em Proceedings of the Tenth Workshop on Statistical Machine
  Translation}, pages 316--322, Lisboa, Portugal.

\bibitem[\protect\citename{Kupiec}1992]{kupiec1992robust}
Julian Kupiec.
\newblock 1992.
\newblock Robust part-of-speech tagging using a hidden markov model.
\newblock {\em Computer Speech \& Language}, 6:225--242.

\bibitem[\protect\citename{M{\`a}rquez and
  Gim{\'e}nez}2004]{marquez2004general}
L~M{\`a}rquez and J~Gim{\'e}nez.
\newblock 2004.
\newblock A general pos tagger generator based on support vector machines.
\newblock {\em Journal of Machine Learning Research}.

\bibitem[\protect\citename{Mikolov \bgroup et al.\egroup
  }2010]{Mikolov:RNNLM:2010}
Tomas Mikolov, Martin Karafiat, Lukas Burget, Jan Cernocky, and Sanjeev
  Khudanpur.
\newblock 2010.
\newblock {Recurrent neural network based language model}.
\newblock In {\em Proceedings of Interspeech}, volume~2, Makuhari, Chiba,
  Japan.

\bibitem[\protect\citename{Mikolov \bgroup et al.\egroup
  }2013a]{Mikolov:MT:2013}
Tomas Mikolov, Quoc~V Le, and Ilya Sutskever.
\newblock 2013a.
\newblock {Exploiting Similarities among Languages for Machine Translation}.
\newblock In {\em CoRR}, pages 1--10.

\bibitem[\protect\citename{Mikolov \bgroup et al.\egroup
  }2013b]{Mikolov:MT1:2013}
Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg~S Corrado, and Jeff Dean.
\newblock 2013b.
\newblock {Distributed representations of words and phrases and their
  compositionality}.
\newblock In {\em Advances in Neural Information Processing Systems}, pages
  3111--3119.

\bibitem[\protect\citename{Mnih and Hinton}2009]{Mnih:2009}
Andriy Mnih and Geoffrey~E. Hinton.
\newblock 2009.
\newblock {A scalable hierarchical distributed language model}.
\newblock In {\em Advances in neural information processing systems}, pages
  1081--1088.

\bibitem[\protect\citename{Morin and Bengio}2005]{Morin:2005}
Frederic Morin and Yoshua Bengio.
\newblock 2005.
\newblock {Hierarchical Probabilistic Neural Network Language Model}.
\newblock In {\em Aistats}, volume~5, pages 246--252.

\bibitem[\protect\citename{Patel and Sasikumar}2016]{patel:rnn-qe:2016}
Raj~Nath Patel and M~Sasikumar.
\newblock 2016.
\newblock {Translation Quality Estimation using Recurrent Neural Network}.
\newblock In {\em Proceedings of the First Conference on Machine Translation},
  volume~2, pages 819--824, Berlin, Germany. Association for Computational
  Linguistics.

\bibitem[\protect\citename{Pimpale and Patel}2015]{pimpale:2015}
Prakash~B. Pimpale and Raj~Nath Patel.
\newblock 2015.
\newblock {Experiments with POS Tagging Code-mixed Indian Social Media Text}.
\newblock {\em ICON}.

\bibitem[\protect\citename{Qin}2015]{qin:pos:2015}
Longlu Qin.
\newblock 2015.
\newblock {POS tagging of Chinese Buddhist texts using Recurrent Neural
  Networks}.
\newblock Technical report, Stanford University.

\bibitem[\protect\citename{Sarkar}2015]{sarkar:2015}
Kamal Sarkar.
\newblock 2015.
\newblock {Part-of-Speech Tagging for Code-mixed Indian Social Media Text at
  ICON 2015}.
\newblock {\em ICON}.

\bibitem[\protect\citename{Schmid and Laws}2008]{schmid2008estimation}
Helmut Schmid and Florian Laws.
\newblock 2008.
\newblock Estimation of conditional probabilities with decision trees and an
  application to fine-grained pos tagging.
\newblock In {\em Proceedings of the 22nd International Conference on
  Computational Linguistics-Volume 1}, pages 777--784. Association for
  Computational Linguistics.

\bibitem[\protect\citename{Schwenk}2007]{Schwenk:2007}
Holger Schwenk.
\newblock 2007.
\newblock {Continuous space language models}.
\newblock In {\em Computer Speech and Language}, volume~21, pages 492--518.

\bibitem[\protect\citename{Socher \bgroup et al.\egroup
  }2013a]{Socher:Parsing:2013}
Richard Socher, John Bauer, Christopher~D. Manning, , and Andrew~Y. Ng.
\newblock 2013a.
\newblock {Parsing With Compositional Vector Grammars}.
\newblock In {\em Proceedings of the ACL 2013}, pages 455--465.

\bibitem[\protect\citename{Socher \bgroup et al.\egroup
  }2013b]{Socher:Sentiment:2013}
Richard Socher, Alex Perelygin, , and Jy~Wu.
\newblock 2013b.
\newblock {Recursive deep models for semantic compositionality over a sentiment
  treebank}.
\newblock In {\em Proceedings of EMNLP}, pages 1631--1642.

\bibitem[\protect\citename{Toutanova \bgroup et al.\egroup
  }2003]{toutanova:2003}
Kristina Toutanova, Dan Klein, Christopher~D Manning, and Yoram Singer.
\newblock 2003.
\newblock {Feature-Rich Part-of-Speech Tagging with a Cyclic Dependency
  Network}.
\newblock In {\em Proceedings of the 2003 Conference of the North American
  Chapter of the Association for Computational Linguistics on Human Language
  Technology-Volume 1}, pages 173--180. Association for Computational
  Linguistics.

\bibitem[\protect\citename{Vyas \bgroup et al.\egroup }2014]{vyas:2014}
Yogarshi Vyas, Spandana Gella, Jatin Sharma, Kalika Bali, and Monojit
  Choudhury.
\newblock 2014.
\newblock Pos tagging of english-hindi code-mixed social media content.
\newblock In {\em EMNLP}, volume~14, pages 974--979.

\bibitem[\protect\citename{Wang \bgroup et al.\egroup }2015]{wang:blstm:2015}
Peilu Wang, Yao Qian, Frank~K Soong, Lei He, and Hai Zhao.
\newblock 2015.
\newblock Part-of-speech tagging with bidirectional long short-term memory
  recurrent neural network.
\newblock {\em arXiv preprint arXiv:1510.06168}.

\bibitem[\protect\citename{Werbos}1990]{Werbos:1990}
Paul~J. Werbos.
\newblock 1990.
\newblock {Backpropagation through time: what it does and how to do it}.
\newblock In {\em IEEE}, volume~78, pages 550--1560.

\bibitem[\protect\citename{Yao \bgroup et al.\egroup }2013]{Yao:2013}
Kaisheng Yao, Geoffrey Zweig, Mei-Yuh Hwang, Yangyang Shi, and Dong Yu.
\newblock 2013.
\newblock {Recurrent neural networks for language understanding}.
\newblock In {\em INTERSPEECH}, pages 2524--2528.

\bibitem[\protect\citename{Yao \bgroup et al.\egroup }2014]{Yao:2014}
Kaisheng Yao, Baolin Peng, Yu~Zhang, Dong Yu, Geoffrey Zweig, and Yangyang Shi.
\newblock 2014.
\newblock {Spoken language understanding using long short-term memory neural
  networks}.
\newblock In {\em Spoken Language Technology Workshop (SLT), IEEE}, pages
  189--194.

\bibitem[\protect\citename{Zeiler}2012]{Zeiler:2002}
Matthew~D. Zeiler.
\newblock 2012.
\newblock {ADADELTA: an adaptive learning rate method}.
\newblock {\em arXiv:1212.5701 [cs.LG]}.

\bibitem[\protect\citename{Zennaki \bgroup et al.\egroup }2015]{zennaki:2015}
Othman Zennaki, Nasredine Semmar, and Laurent Besacier.
\newblock 2015.
\newblock {Unsupervised and Lightly Supervised Part-of-Speech Tagging Using
  Recurrent Neural Networks}.
\newblock In {\em Proceedings of the 29th Pacific Asia Conference on Language,
  Information and Computation}, Shanghai, China.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{eacl2017}
\usepackage{times}
\usepackage[hyphens]{url}
\usepackage{latexsym}
\usepackage{color}
\usepackage{comment}
\usepackage{xspace}
\usepackage{mathtools}

\pdfoutput=1

\newcommand{\note}[1]{\textit{\small\color{blue}{#1}}}

\newcommand{\hien}{\textit{hi-en}\xspace}
\newcommand{\bnen}{\textit{bn-en}\xspace}
\newcommand{\teen}{\textit{te-en}\xspace}

\newcommand{\en}{\textit{en}\xspace}
\newcommand{\hi}{\textit{hi}\xspace}

\newcommand{\fb}{\textit{fb}\xspace}
\newcommand{\twt}{\textit{twt}\xspace}
\newcommand{\wa}{\textit{wa}\xspace}

\newcommand{\rnnlm}{\textit{RNN-LM}\xspace}
\newcommand{\rnn}{RNN\xspace}
\newcommand{\srnn}{\textit{Simple\_RNN}\xspace}
\newcommand{\lstm}{\textit{LSTM}\xspace}
\newcommand{\dlstm}{\textit{Deep LSTM}\xspace}
\newcommand{\gru}{\textit{GRU}\xspace}
\newcommand{\grup}{\textit{GRU\_Pre}\xspace}
\newcommand{\grupl}{\textit{GRU\_Pre\_Lang}\xspace}

\newcommand{\stan}{\textit{Stanford}\xspace}
\newcommand{\hun}{\textit{HunPos}\xspace}

\eaclfinalcopy 

\title{Recurrent Neural Network based Part-of-Speech Tagger for Code-Mixed Social Media Text}

\author{Raj Nath Patel \\
	KBCS, CDAC Mumbai \\
	{\tt rajnathp@cdac.in} \\\And
	Prakash B. Pimpale \\
	KBCS, CDAC Mumbai \\
	{\tt prakash@cdac.in} \\\And
	Sasikumar M. \\
	KBCS, CDAC Mumbai \\
	{\tt sasi@cdac.in} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
This paper describes Centre for Development of Advanced Computing's (CDACM) submission to the shared task-'Tool Contest on POS tagging for Code-Mixed Indian Social Media (Facebook, Twitter, and Whatsapp) Text', collocated with ICON-2016. The shared task was to predict Part of Speech (POS) tag at word level for a given text. The code-mixed text is generated mostly on social media by multilingual users. The presence of the multilingual words, transliterations, and spelling variations make such content linguistically complex. In this paper, we propose an approach to POS tag code-mixed social media text using Recurrent Neural Network Language Model (\rnnlm) architecture. We submitted the results for Hindi-English (\hien), Bengali-English (\bnen), and Telugu-English (\teen) code-mixed data.
\end{abstract}

\section{Introduction}

Code-Mixing and Code-Switching are observed in the text or speech produced by a multilingual user. Code-Mixing occurs when a user changes the language within a sentence, i.e. a clause, phrase or word of one language is used within an utterance of another language. Whereas, the co-occurrence of speech extract of two different grammatical systems is known as Code-Switching.

The language analysis of code-mixed text is a non-trivial task. Traditional approaches of POS tagging are not effective, for this text, as it does not adhere to any grammatical structure in general. Many studies have shown that \rnn based POS taggers produced comparable results and, is also the state-of-the-art for some languages. However, to the best of our knowledge, no study has been done for \rnn based POS tagging of code-mixed data. 

In this paper, we have proposed a POS tagger using \rnnlm architecture for code-mixed Indian social media text. Earlier, researchers have adopted \rnnlm architecture for Natural language Understanding (NLU)~\cite{Yao:2013,Yao:2014} and Translation Quality Estimation~\cite{patel:rnn-qe:2016}. \rnnlm models are similar to other vector-space language models~\cite{Bengio:2003,Morin:2005,Schwenk:2007,Mnih:2009} where we represent each word with a high dimensional real-valued vector. We modified \rnnlm architecture to predict the POS tag of a word, given the word and its context. Let's consider the following example: \\

Input: $behen\ ki\ shaadi\ and\ m\ not\ there$

Output : G\_N G\_PRP G\_N CC G\_V G\_R G\_R \\

In the above sentence, to predict POS tag (G\_N) for the word ‘$shaadi$’ using an \rnnlm model with window size 3, the input will be ‘$ki\ shaadi\ and$’. Whereas, in standard \rnnlm model, ‘$ki\ and$’ will be the input with ‘$shaadi$’ as the output. We will discuss details of various models tried and their implementations in section~\ref{settings}.

In this paper, we show that our approach achieves results close to the state-of-the-art systems such as \footnotemark \stan~\cite{toutanova:2003}, \footnotetext{\url{http://nlp.stanford.edu/software/tagger.shtml} (Maximum-Entropy based POS tagger)} and \footnotemark \hun~\cite{hunpos:2007} \footnotetext{\url{https://code.google.com/archive/p/hunpos/} (Hidden Markov Model based POS tagger)}.

\section{Related Work}

POS tagging has been investigated for decades in the literature of Natural Language Processing (NLP). Different methods like a Support Vector Machine~\cite{marquez2004general}, Decision Tree~\cite{schmid2008estimation}, Hidden Markov Model (HMM)~\cite{kupiec1992robust} and, Conditional Random Field Auto Encoders~\cite{ammar2014conditional} have been tried for this task. Among these works, Neural Network (NN) based models is mainly related to this paper. In NN family, \rnn is widely used network for various NLP applications~\cite{Mikolov:RNNLM:2010,Mikolov:MT:2013,Mikolov:MT1:2013,Socher:Parsing:2013,Socher:Sentiment:2013}.

Recently, \rnn based models have been used to POS tag the formal text, but have not been tried yet on code-mixed data. \newcite{wang:blstm:2015} have tried Bidirectional Long Short-Term Memory (LSTM) on Penn Treebank WSJ test set, and reported state-of-the-art performance. \newcite{qin:pos:2015} has shown that \rnn models outperform Majority Voting (MV) and HMM techniques for POS tagging of Chinese Buddhist text. \newcite{zennaki:2015} have used \rnn for resource-poor languages and reported comparable results with state-of-the-art systems ~\cite{das:2011,duong:2013,gouws:2015}.

Work on POS tagging code-mixed Indian social media text is at a very nascent stage to date. \newcite{vyas:2014} and \newcite{jamatia:2015} have worked on data labeling and automatic POS tagging of such data using various machine learning techniques. Building further on that labeled data, \newcite{pimpale:2015} and, \newcite{sarkar:2015} have tried word embedding as an additional feature to the machine learning based classifiers for POS tagging.


\section{Experimental Setup}
\label{settings}
\subsection{RNN Models}
There are many variants of \rnn networks for different applications. For this task, we used elaman~\cite{elman:1990}, Long Short-Term Memory (\lstm)~\cite{Hochreiter:1997}, \dlstm, Gated Recurrent Unit (\gru) ~\cite{Cho:2014}, which are widely used \rnn models in the NLP literature.

In the following sub-sections, we gave a brief description of each model with mathematical equations (1,2, and 3). In the equations, $x_t$ and $y_t$ are the input and output vectors respectively. $h_t$ and $h_{t-1}$ represent the current and previous hidden states respectively. $W_*$ are the weight matrices and $b_*$ are the bias vectors. $\odot$ is the elementwise multiplication of the vectors. We used $sigm$, the logistic sigmoid and $tanh$, the hyperbolic tangent function to add nonlinearity in the network with $softmax$ function at the output layer. 

\subsubsection{ELMAN}
Elman and Jordon~\cite{jordan:1986} networks are the simplest network in \rnn family and are known as \srnn. Elman network is defined by the following set of equations:
\begin{flalign}
& h_t = sigm(W_{xh}x_t + W_{hh}h_{t-1} + b_h) \\
& y_t = softmax(W_{hy}h_t + b_y) \notag
\end{flalign}

\subsubsection{LSTM}
\lstm is found to be better for modeling of long-range dependencies than \srnn. \srnn also suffers from the problem of vanishing and exploding gradient  ~\cite{Yoshua:1994}. \lstm and other complex \rnn models tackle this problem by introducing a gating mechanism. Many variants of \lstm~\cite{Graves:2013,Yao:2014,Jozefowicz:2015} have been tried in literature for the various tasks. We implemented the following version:
\begin{flalign}
& i_t = sigm(W_{xi}x_t + W_{hi}h_{t-1} + b_i) \\
& o_t = sigm(W_{xo}x_t + W_{ho}h_{t-1} + b_o) \notag \\
& f_t = sigm(W_{xf}x_t + W_{hf}h_{t-1} + b_f) \notag \\
& j_t = tanh(W_{xj}x_t + W_{hj}h_{t-1} + b_j) \notag \\
& c_t = c_{t-1} \odot f_t + i_t \odot j_t  \notag \\
& h_t = tanh(c_t) \odot o_t \notag \\
& y_t = softmax(W_{hy}h_t + b_y) \notag
\end{flalign}
where $i$, $o$, $f$ are $input$, $output$ and $forget$ gates respectively. $j$ is the new memory content and $c$ is updated memory. 

\subsubsection{Deep LSTM}
In this paper, we used \dlstm with two layers. \dlstm is created by stacking multiple \lstm on the top of each other. The output of lower \lstm forms input to the upper \lstm. For example, if $h_t$ is the output of lower \lstm, then we apply a matrix transform to form the input $x_t$ for the upper \lstm. The Matrix transformation enables us to have two consecutive \lstm layers of different sizes.

\subsubsection{GRU}
\gru is quite a similar network to the \lstm, without any memory unit. \gru network also uses a different gating mechanism with $reset$ ($r$) and $update$ ($z$) gates. The following set of equations defines a \gru model:
\begin{flalign}
& r_t = sigm(W_{xr}x_t + W_{hr}h_{t-1} + b_r) \\
& z_t = sigm(W_{xz}x_t + W_{hz}h_{t-1} + b_z) \notag \\
& \widetilde{h}_t = tanh(W_{xh}x_t + W_{hh}(r_t \odot h_{t-1}) + b_h) \notag \\
& h_t = z_t \odot h_{t-1} + (1 - z_t) \odot \widetilde{h}_t \notag \\
& y_t = softmax(W_{hy}h_t + b_y) \notag
\end{flalign}

\subsection{Implementation}
\label{subsec:impl}
All the models were implemented using \footnotemark THEANO \footnotetext{\url{http://deeplearning.net/software/theano/\#download}} framework~\cite{Bergstra:2010,Bastien:2012}. For all the models, the word embedding dimensionality was 100, no of hidden units were 100 and the context word window size was 5 ($w_{i-2}w_{i-1}w_iw_{i+1}w_{i+2}$). We initialized all the square weight matrices as random orthogonal matrices. All the bias vectors were initialized to zero. Other weight matrices were sampled from a Gaussian distribution with mean 0 and variance $0.0001$.

We trained all the models using Truncated Back-Propagation-Through-Time (T-BPTT)~\cite{Werbos:1990} with the stochastic gradient descent. Standard values of hyper-parameters  were used for RNN model training, as suggested in the literature~\cite{Yao:2014,patel:rnn-qe:2016}. The depth of BPTT was fixed to 7 for all the models. We trained each model for 50 epochs and used Ada-delta~\cite{Zeiler:2002} to adapt the learning rate of each parameter automatically ($\epsilon = 10^{-6}$ and $\rho = 0.95$).

\subsection{Data} \label{sec:data}
We used the data shared by the contest organizers~\cite{jamatia:2016}. The code-mixed data of \bnen, \hien and \teen was shared separately for the Facebook (\fb), Twitter (\twt) and Whatsapp (\wa) posts and conversations with Coarse-Grained (CG) and Fine-Grained (FG) POS annotations. We combined the data from \fb, \twt, and \wa for CG and FG annotation of each language pair. The data was divided into training, testing, and development sets. Testing and development sets were randomly sampled from the complete data. Table~\ref{tab:data} details sizes of the different sets at the sentence and token level. Tag-set counts for CG and FG are also provided.

We preprocess the text for Mentions, Hashtags, Smilies, URLs, Numbers and, Punctuations. In the preprocessing, we mapped all the words of a group to a single new token as they have the same POS tag. For example, all the Mentions like @dhoni, @bcci, and @iitb were mapped to @user; all the Hashtags like \#dhoni, \#bcci, \#iitb were mapped to \#user.

\begin{table*}[hbt]
	\centering
	\begin{tabular}{l|ccc|ccc|cc}
		\multicolumn{1}{c|}{} & \multicolumn{3}{c|}{\#sentences} & \multicolumn{3}{c|}{\#tokens} & \multicolumn{2}{c}{\#tags} \\
		\multicolumn{1}{l|}{code-mix} & \multicolumn{1}{l}{training} & \multicolumn{1}{l}{dev} & \multicolumn{1}{l|}{testing} &  \multicolumn{1}{l}{training} & \multicolumn{1}{l}{dev} & \multicolumn{1}{l|}{testing} &  \multicolumn{1}{l}{CG} & \multicolumn{1}{l}{FG} \\ \hline
		\multicolumn{1}{c|}{\hien} & 2430 & 100 & 100 & 37799 & 1888 & 1457 & 18 & 40 \\
		\multicolumn{1}{c|}{\bnen} & 524 & 50 & 50 & 11977 & 1477 & 1231 & 18 & 38 \\
		\multicolumn{1}{c|}{\teen} & 1779 & 100 & 100 & 26470 & 1436 & 1543 & 18 & 50 \\
	\end{tabular}
	\caption{Data Distribution; CG: Coarse-Grained, FG: Fine-Grained}
	\label{tab:data}
\end{table*}

\subsection{Methodology} \label{sec:methodology}
The \rnnlm models use only the context words' embedding as the input features. We experimented with three \rnn model configurations. In the first setting (\srnn, \lstm, \dlstm, \gru), we learn the word representation from scratch with the other model parameters. In the second configuration (\grup), we trained word representations (pre-training) using $word2vec$~\cite{Mikolov:MT1:2013} tool and fine tuned with the training of other parameters of the network. Pre-training not only guides the learning towards minima with better generalization in non-convex optimization~\cite{bengio:2009,erhan:2010} but also improves the accuracy of the system~\cite{kreutzer:2015,patel:rnn-qe:2016}. In the third setting (\grupl), we also added language of the words as an additional feature with the context words. We learn the vector representation of languages similar to that of words, from scratch.

\section{Results}
We used F1-Score to evaluate the experiments, results are displayed in the Table~\ref{tab:results}. We trained models as described in the section~\ref{sec:methodology}. To compare our results, we also trained the \stan and \hun taggers on the same data, accuracy is given in Table~\ref{tab:results}.

From the table, it is evident that pre-training and language as an additional feature is helpful. Also, the accuracy of our best system (\grupl) is comparable to that of \stan and \hun. \gru models are out-performing other models (\srnn, \lstm, \dlstm) for this task also as reported by~\newcite{Chung:2014} for a suit of NLP tasks.

\begin{table*}[hbt]
	\centering
	\begin{tabular}{l|cc|cc|cc}
		\multicolumn{1}{l|}{} & \multicolumn{2}{c|}{\hien \%F1 score} & \multicolumn{2}{c|}{\bnen \%F1 score} & \multicolumn{2}{c}{\teen \%F1 score} \\
		\multicolumn{1}{l|}{model} & \multicolumn{1}{c}{CG} & \multicolumn{1}{c|}{FG} & \multicolumn{1}{c}{CG} & \multicolumn{1}{c|}{FG} & \multicolumn{1}{c}{CG} & \multicolumn{1}{c}{FG} \\ \hline
		\multicolumn{1}{l|}{\srnn} & 78.16 & 68.73 & 70.16 & 64.49 & 72.27 & \bf 69.04 \\
		\multicolumn{1}{l|}{\lstm} & 62.75 & 53.94 & 41.91 & 35.05 & 57.59 & 51.45 \\
		\multicolumn{1}{l|}{\dlstm} & 70.07 & 59.78 & 54.64 & 46.88 & 65.86 & 59.45 \\
		\multicolumn{1}{l|}{\gru} & \bf 78.29 & \bf 69.32 & \bf 71.90 & \bf 64.96 & \bf 72.40 & 68.72 \\ \hline
		\multicolumn{1}{l|}{\grup} & 80.51 & 71.72 & 74.77 & 68.54 & 74.02 & 70.05 \\
		\multicolumn{1}{l|}{\grupl} & 80.92 & 73.10 & 74.05 & 69.23 & 74.00 & 70.33 \\ \hline
		\multicolumn{1}{l|}{\hun} & 77.50 & 69.04 & 76.55 & 71.02 & 74.30 & 70.73 \\
		\multicolumn{1}{l|}{\stan} & 79.89 & 73.91 & 79.36 & 73.44 & 77.05 & 73.42
	\end{tabular}
	\caption{F1 scores for different experiments}
	\label{tab:results}
\end{table*}

\section{Submission to the Shared Task} \label{sec:submission}
The contest was having two type of submissions, first, $constrained$: restricted to use only the data shared by the organizers with the participants' implemented systems; second, $unconstrained$: participants were allowed to use the publicly available resources (training data, implemented systems etc.).

We submitted for all the language pairs (\hien, \bnen and, \teen) and domains (\fb, \twt and, \wa). For constrained submission, the output of \grupl was used. We trained \stan POS tagger with the same data for $unconstrained$ submission.~\newcite{jamatia:2016} evaluated all the submitted systems against another gold-test set and reported the results.



\section{Analysis}
We did a preliminary analysis of our systems and reported few points in this section. 
\begin{itemize}
	\item The POS categories, contributing more in the error are G\_X, G\_V, G\_N and G\_J for coarse-grained and V\_VM, JJ, N\_NN and N\_NNP for fine-grained systems. Also, we did the confusion matrix analysis and found that these POS tags are mostly confused with each other only. For instance, G\_J POS tag was tagged 28 times wrongly to the other POS tags in which 17 times it was G\_N. 
	
	\item \rnn models require a huge amount of corpus to train the model parameters. From the results, we can observe that for \hien and \teen with only approx 2K training sentences, the results of best \rnn model (\grupl) are comparable to \stan and \hun. For \bnen, the corpus was very less (only approx 0.5K sentences) for \rnn training which resulted into poor performance compared to \stan and \hun. With this and the earlier work on RNN based POS tagging, we can expect that \rnn models could achieve state-of-the-art accuracy with given the sufficient amount of training data. 
	
	\item In general, \lstm and \dlstm models perform better than \srnn. But here, \srnn is outperforming both \lstm and \dlstm. The reason could be less amount of data for training such a complex model. 
	
	\item Few orthographically similar words of English and Hindi, having different POS tags are given with examples in Table~\ref{tab:word-list}. System confuses in POS tagging of such words. With adding language as an additional feature, we were able to tag these type of words correctly.
	
\end{itemize}

\begin{table}[!hbt]
	\centering
	\begin{tabular}{l|c|l|c}
		word & lang &  example & POS \\ \hline
		are & \hi & \textbf{are} shyaam kidhar ho? & PSP \\
		are & \en & they \textbf{are} going. & G\_V \\ \hline
		to & \hi & tumane \textbf{to} dekha hi nhi. & G\_PRT \\
		to & \en & they go \textbf{to} school. & CC \\ \hline
		hi & \hi & mummy to aisi \textbf{hi} hain. & G\_V \\
		hi & \en & \textbf{hi}, how are you. & G\_PRT
	\end{tabular}
	\caption{Similar words in \hien data}
	\label{tab:word-list}
\end{table}

\section{Conclusion and Future Work}
We developed language independent and generic POS tagger for social media text using \rnn networks. We tried \srnn, \lstm, \dlstm and, \gru models. We showed that \gru outperforms other models, and also benefits from pre-training and language as an additional feature. Also, the accuracy of our approach is comparable to that of \stan and \hun.

In the future, we could try \rnn models with more features like POS tags of context words, prefixes and suffixes, length, position, etc. Word characters also have been found to be a very useful feature in \rnn based POS taggers.

\bibliography{eacl2017}
\bibliographystyle{eacl2017}

\end{document}